Thinking and Inquiry Two train whistles, A and B each have a frequency of 392Hz. A is at rest while B is moving at 35m/s [right]. A railway worker moving at 15 m/s[left] is between the two trains. Air temp is 293 K. 14. a) What is the frequency heard by the railway worker from Train A? (3)
b) What is the frequency heard by the railway worker by train B? (3) c) What is the beat frequency heard by the railway worker caused by the two trains? (2)
Answer #1

Part A.

Using Doppler's effect:

When Observer is moving away from Stationary Source

f1 = f0*(V - Vo)/V

Vo = Speed of Observer = 15 m/sec

V = Speed of Sound at 293 K = 343 m/sec

f0 = frequency of Source = 392 Hz

So,

f1 = frequency heard by worker by train A = ?

f1 = 392*(343 - 15)/343

f1 = 374.8 Hz

Part B.

Since train is moving in right and worker is moving in left, So

When Source and observer both moving away from each other, then

f2 = f0*(V - Vo)/(V + Vs)

Vs = Speed of source = 35 m/sec

Using these values:

f2 = 392*(343 - 15)/(343 + 35)

f2 = 340.1 Hz = frequency heard by worker by train B

Part C.

Beat frequency = f1 - f2

Beat frequency = 374.8 - 340.1 = 34.7 Hz

Please Upvote.